|02/28/2001||Tuesday February 27, ~10:40 - 11:00 - The web server, webgrp1, experienced a series of system errors taking it and all hosted web sites offline for approximately 20 minutes. Service has since been restored. This was a repeat occurrence of the problem reported on 2/6/01. A solution has been found and this problem is not expected to occur again.|
|03/17/2001||Wednesday March 14, ~9:15 am - Web services were briefly interrupted on the webgrp1 server. Service was restored to all hosted websites by ~9:25 am. This interruption initially appears to be related to possible problem with PERL service. More information will be posted as this problem is further documented.|
|03/17/2001||Wednesday March 14 - Some areas of Dobson Hall lost network connectivity from ~7pm until 8:30am. This outage was caused by an electrical service failure. Physical Plant electricians have corrected the problem.|
|03/17/2001||Thursday March 15 - The IBM Mainframe crashed at ~ 2:45 PM due to a channel error. This effected some internal library operations. Service was restored at ~ 3:30 PM.|
|03/29/2001||Monday evening March 26 - The Groups web server Argon was down for an undetermined amount of time between 7:15 and 9:45. After experiencing a series of internal service errors, Argon was restarted at 9:53pm . Service to all hosted web sites was restored by 10pm.|
|04/12/2001||On Sunday April 8 some residents in Blanton Hall lost network service from
about 3 p.m. until 6 p.m. The problem was caused by an electrical power problem.|
|04/15/2001||Thursday April 12 - The Pickler Library network is down. A core network backbone switch was damaged by a electrical power problem. Repairs are currently underway.
Thursday April 12 - The Pickler Library network was down from ~ 7:00 AM until 10:00 AM. A core network backbone switch was damaged by a electrical power problem.
Thursday April 12 - Network service to Missouri Hall was down from ~ 7:30 AM until 9:05 AM due to an electrical power problem
|05/17/2001||The Groups Web Server, Argon, has been restarted several times during the past two weeks as updates and maintenance packs were applied.
Any service outages should have been brief.
Webmasters may, however, need to reapply custom file permissions.
|06/14/2001||The Groups Web server, Argon, experienced a hardware failure Monday evening (6/11). Services have been restored as of 9:15am Tuesday (6/12). As additional checks and repairs are performed, brief outages may occur Tuesday and Wednesday.|
|06/29/2001||Due to a major electrical power outage on the Truman power grid, many network/web services were interrupted or unavailable from Sunday evening, 6/24 through Tuesday 6/26. All services were restored by 2a.m. Wednesday morning.
|07/01/2001||The Gold server was taken off-line for scheduled upgrades and maintenance at 5pm Friday (6/29). Service was restored by 9pm. During this time frame, e-mail, web sites with www2. addresses and all other services supported on Gold were unavailable. Campus dial-up connections were also be unavailable during this outage.|
|07/04/2001||The Argon server was taken off-line for scheduled maintenance at 5:30 pm Sunday, 7/1. Service was restored by 6pm. All hosted web sites were unavailable for a brief amount of time as the server was rebooted.|
|07/05/2001||The Argon server went down last night (Monday, 7/2) at an undetermined time between 7pm and 11:00pm. Service was restored by 8am Wednesday morning. The cause of the failure is being investigated at this time.|
|07/22/2001||oth the DK HelpDesk system server (TS11) and the Group Web Server (Argon) experienced a variety of technical difficulties on Thursday, 7/20/01. Some pages were unavailable and Web Services were interrupted a number of times. Behavior indicates the problems may have been caused by complications from an attack of the "Code Red" Worm virus. Both servers have had appropriate service packs and security updates applied and service has been restored as of 9:30am. |
|08/19/2001||On Sunday, August 19th, 2001, from 8:00 am - 12:00 pm, MOREnet will perform maintenance on core equipment in the Kansas City area. This maintenance is to replace an equipment failure identified within the Kansas City ATM switch.
MOREnet customers in the Kansas City area will experience a loss of connectivity during this maintenance period as the hardware swap is performed.
For questions or comments concerning this announcement, please reply to: email@example.com|
|09/10/2001||On Friday evening, 9/7/01, at approximately 11:00pm, a power outage in McClain Hall took the Truman Data Center offline. Power was later restored and Internet access and most other network services were back online by 9am Saturday morning. The web server Krypton was not fully operational till 12noon on Monday.|
|09/22/2001||On Wednesday, September 19, there was a partial disruption of network services for some Truman network users from 8 AM until 9:30 AM. This problem was caused by a Truman user running an unauthorized and mis-configured DHCP server.|
|09/29/2001||Internet access was unavailable on Thursday, September 27 from 7:30 AM until 7:55 AM. This was due to MOREnet performing emergency unscheduled maintenance on the state network. |
|09/30/2001||Maintenance scheduled for Sunday, September 30, 2001.
MOREnet will be conducting maintenance on its Jefferson City BPX at 6:00 a.m., September 30th. The maintenance period is not expected to exceed two hours; Customers will experience a brief loss of connectivity during this maintenance period.|
|10/06/2001||Web Services on Argon (www) were unavailable Wednesday, October 3 from approximately 6pm to 7:15pm. Service was restored after the system was rebooted.|
|10/07/2001||The main web server ARGON will be taken offline at 9:00PM on Monday,
October 8, for maintenance. Unless unforeseen problems arise, web services will be restored by 8:00AM on Tuesday morning. Many of Truman's websites will not be available during this time.|
|10/05/2001||The email and personal website server, GOLD, will be taken offline today (Friday, October 5th) around 6:00PM for upgrades and maintenance.
Unless unforeseen problems arise, Gold will be back online and available
to users no later than 8:00PM. Email and personal websites will not be available during this time.|
|10/14/2001||The email and personal web server, Gold, was rebooted this morning at 7:50 after maintenance to resolve incorrect date stamps and other issues affecting e-mail services. Service was restored by 8:15.|
|10/18/2001||On Tuesday, October 16 at ~10:14 AM some primary DNS services were disrupted due to a problem with the server. All primary DNS services were restored shortly before 10:20 AM. The secondary DNS server remained up and active during this brief period.|
|10/20/2001||On Thursday evening, October 18 at ~5:00pm the administrative web server Argon was taken offline for maintenance for approximately 30 minutes. Most hosted web sites were back up at 5:30. The Truman home site was off-line from ~5:30 - 6:30pm. All website administrators were notified prior to the maintenance window.